OpenStudy (anonymous):

If 10^100 googol, then 1000^100 equals...

OpenStudy (anonymous):

what is googol?

OpenStudy (anonymous):

it's just an expression thing in this question (:

OpenStudy (anonymous):

1000^100 = (10^3)^100 = (10^100)^3 = googol^3

OpenStudy (anonymous):

wow math notation, today i learn a new one....

OpenStudy (anonymous):

haha, but i'm not really sure myself though >,>

OpenStudy (anonymous):

1 googol = 1.0 × 10^100

OpenStudy (anonymous):

Thanks everyone :3

OpenStudy (anonymous):

If it weren't for you guys i would have guessed 100googol as my answer :L

OpenStudy (anonymous):

googol is just a number unit, like a thousand or a hundred 10^2 = hundred 10^3 = thousand 10^6 = million 10^9 = billion 10^100 = googol

OpenStudy (anonymous):

I understand now :3

OpenStudy (anonymous):

the search engine is named after it

OpenStudy (anonymous):

OMG it is! :D But it's just a bit different

OpenStudy (anonymous):

that's why it's like Goooooooooooooooooooooogle at the bottom a reference to the number

OpenStudy (anonymous):

because one googol has one hundred zeros

OpenStudy (anonymous):

seriously...i thought she mentioned about Google in the beginning hahaha

OpenStudy (anonymous):

:D LOLS I learnt like so much in the past 15 minutes already :)

OpenStudy (anonymous):

learning sucks

OpenStudy (anonymous):

no it does not D:

OpenStudy (anonymous):

seriously i hate learning....but if i dont learn i can do nothing :P

OpenStudy (anonymous):

Exactly. Okay goodnight guys :)

OpenStudy (anonymous):

i'll googol it! (sorry guys!)